The treasurer of a municipality claims that the average net worth of families living in this municipality is ₱590,000. A random sample of 50 families selected from this area produced a mean net worth of ₱720,000 with standard deviation of ₱65,000. Using 1% significance level, can we conclude that the claim is true? Also, find the 99% confidence interval of the true mean.


1
Expert's answer
2022-05-09T16:20:28-0400

We need to construct the "99\%" confidence interval for the population mean "\mu."

The critical value for "\alpha = 0.05" and "df = n-1 =49" degrees of freedom is "t_c = z_{1-\alpha/2; n-1} = 2.679952."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-t_c\times\dfrac{s}{\sqrt{n}},\bar{x}+t_c\times\dfrac{s}{\sqrt{n}})"

"=(720000-2.679952\times\dfrac{65000}{\sqrt{50}},"

"720000+2.679952\times\dfrac{65000}{\sqrt{50}})"

"=(695364.841,744635.159)"

Therefore, based on the data provided, the 99% confidence interval for the population mean is "695364.841 < \mu < 744635.159," which indicates that we are 99% confident that the true population mean "\mu" is contained by the interval "(695364.841, 744635.159)."


The following null and alternative hypotheses need to be tested:

"H_0:\mu=590000"

"H_a:\mu\not=590000"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.01," "df=n-1=49" degrees of freedom, and the critical value for a two-tailed test is "t_c =2.679952."

The rejection region for this two-tailed test is "R = \{t: |t|> 2.679952\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{720000-590000}{65000/\sqrt{50}}\approx14.142"

Since it is observed that "|t |= 14.142 \ge2.679952=t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for two-tailed test, "df=49" degrees of freedom, "t=14.142" is "p = 0," and since "p = 0< 0.01=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is different than 590000, at the "\alpha = 0.01" significance level.